how many circles of radius 7cm can be cut from the paper of length 50cm and width 32cm
Answer by thejackal(72)

area of paper = L * W = 50*32 = 1600cm^2;
area of a circle = pi * r^2 = 22/7 * (7*7) = 22/7*49 = 22 * 7 = 154cm^2
surface area of paper/ surface area of circle = 1600/154 = 800/77 or 10.3896
but since we are only interested in complete circles, the answer is 10.